# Dynamic Gaussian Splatting from Defocused and Motion-blurred Monocular Videos

### v1

@@ Abstract
This paper presents a unified framework that allows high-quality dynamic Gaussian Splatting from both defocused and motion-blurred monocular videos. Due to the significant difference between the formation processes of defocus blur and motion blur, existing methods are tailored for either one of them, lacking the ability to simultaneously deal with both of them. Although the two can be jointly modeled as blur kernel-based convolution, the inherent difficulty in estimating accurate blur kernels greatly limits the progress in this direction. In this work, we go a step further towards this direction. Particularly, we propose to estimate per-pixel reliable blur kernels using a blur prediction network that exploits blur-related scene and camera information and is subject to a blur-aware sparsity constraint. Besides, we introduce a dynamic Gaussian densification strategy to mitigate the lack of Gaussians for incomplete regions, and boost the performance of novel view synthesis by incorporating unseen view information to constrain scene optimization. Extensive experiments show that our method outperforms the state-of-the-art methods in generating photorealistic novel view synthesis from defocused and motion-blurred monocular videos. Our code and trained model will be made publicly available.
@@ 2 Related Works
Image deblurring. Early works on image and video deblurring [21, 22, 51, 38, 67] primarily rely on CNNs [48] and RNNs [71]. Later approaches enhance the deblurring process by incorporating additional cues such as depth maps [41, 27, 52, 1], light fields [45, 46], and 3D geometry [39, 40, 57, 61]. Leveraging multi-view and 3D information, NeRF- and 3DGS-based deblurring methods can generate sharp images with improved scene consistency [43, 23, 36, 24, 25, 9]. Among these, some approaches [36, 24, 35, 3] model blur by predicting deformable sparse kernels with per-pixel bases. Others simulate the physical formation of blur by integrating multiple sharp images over the exposure time [70, 53, 40] or use a camera model that learns aperture and focal length parameters [57, 61]. Additionally, some works [49, 34, 60, 32, 4] incorporate 3D geometry into monocular video deblurring. However, existing methods are typically designed for either defocus blur or motion blur. In contrast, our approach effectively removes both motion blur and defocus blur from monocular videos, producing high-quality novel-view images of dynamic scenes.
@@ 3.2 Unified Blur Synthesis
Figure 3: Visual comparison of novel view synthesis on the defocus blur dataset from [34].
@@ 3.3 Training Strategy
Figure 4: Visual comparison of novel view synthesis on the defocus blur dataset from [34]. Here, we also compare with methods fed with deblurred images produced by a state-of-the-art video deblurring method [68] to manifest the effectiveness of our method.
@@ 3.4 Loss Function
Figure 5: Visual comparison of novel view synthesis on the motion blur dataset from [49].
@@ B.1 Novel View Synthesis Comparison
D2RF-v2 and DyBluRF-v2 Dataset. In the Table 6, we evaluate our method on two datasets (DyBluRF-v2 and D2RF-v2) with both motion and defocus blur occurring simultaneously. Note, we obtain the DyBluRF-v2 dataset by applying depth-of-field (DoF) rendering technique in Bokehme [44] to the original DyBluRF dataset [49] to simulate defocus blur, and create the D2RF-v2 dataset with motion blur by processing the original D2RF dataset [34] using the motion blur generation method in Davanet [72]. As shown, our method clearly outperforms all the compared methods on the two datasets, verifying the advantage of our method in handling cases with motion and defocus blur occurring jointly.
B.2 Deblurring Comparison
Appendix C Additional Ablation Results
C.1 Ablation on Blur Kernel Size
Table 7 further perform quantitative evaluation on how different blur kernel sizes (denoted as $K$) affect the performance of our method. As shown, a larger blur kernel helps to obtain better results. However, this trend becomes less obvious when $K$ is larger than 9. To balance the performance and the computational cost, we thus choose KK=9 as our default choice.
Appendix D Robustness to Preprocessing and Segmentation Errors
@@
Figure 10: Visual comparison of novel view synthesis on the motion blur dataset from [49]. Here, we also compare with methods fed with deblurred images produced by a state-of-the-art video deblurring method [68] to manifest the effectiveness of our method.
@@
Table 6: Quantitative comparison of novel view synthesis on the D2RF-v2 defocus blur dataset and the DyBluRF-v2 motion blur dataset.
@@
D3DGS [65] 23.66 0.739 0.257 21.75 0.655 0.289
SoM [54] 29.04 0.820 0.094 27.28 0.791 0.103
D2RF [34] 27.82 0.795 0.132 25.89 0.722 0.133
DyBluRF [49] 27.30 0.771 0.150 26.54 0.753 0.112
De4DGS [60] 29.74 0.856 0.078 27.97 0.824 0.087
Ours 30.26 0.885 0.062 28.55 0.859 0.064
Table 7: Effect of varying blur kernel size $K$.

### v3

@@ Abstract
This paper presents a unified framework that allows high-quality dynamic Gaussian Splatting from both defocused and motion-blurred monocular videos. Due to the significant difference between the formation processes of defocus blur and motion blur, existing methods are tailored for either one of them, lacking the ability to simultaneously deal with both of them. Although the two can be jointly modeled as blur kernel-based convolution, the inherent difficulty in estimating accurate blur kernels greatly limits the progress in this direction. In this work, we go a step further towards this direction. Particularly, we propose to estimate per-pixel reliable blur kernels using a blur prediction network that exploits blur-related scene and camera information and is subject to a blur-aware sparsity constraint. Besides, we introduce a dynamic Gaussian densification strategy to mitigate the lack of Gaussians for incomplete regions, and boost the performance of novel view synthesis by incorporating unseen view information to constrain scene optimization. Extensive experiments show that our method outperforms the state-of-the-art methods in generating photorealistic novel view synthesis from defocused and motion-blurred monocular videos. Our code is available at https://github.com/hhhddddddd/dydeblur.
@@ 2 Related Works
Image deblurring. Early works on image and video deblurring [22, 23, 52, 39, 69] primarily rely on CNNs [49] and RNNs [73]. Later approaches enhance the deblurring process by incorporating additional cues such as depth maps [42, 28, 53, 1], light fields [46, 47], and 3D geometry [40, 41, 58, 62, 64]. Leveraging multi-view and 3D information, NeRF- and 3DGS-based deblurring methods can generate sharp images with improved scene consistency [44, 24, 37, 25, 26, 10]. Among these, some approaches [37, 25, 36, 3, 6] model blur by predicting deformable sparse kernels with per-pixel bases. Others simulate the physical formation of blur by integrating multiple sharp images over the exposure time [72, 54, 41] or use a camera model that learns aperture and focal length parameters [58, 62]. Additionally, some works [50, 35, 61, 33, 4] incorporate 3D geometry into monocular video deblurring. However, existing methods are typically designed for either defocus blur or motion blur. In contrast, our approach effectively removes both motion blur and defocus blur from monocular videos, producing high-quality novel-view images of dynamic scenes.
@@ 3.2 Unified Blur Synthesis
Figure 3: Visual comparison of novel view synthesis on the D2RF defocus blur dataset [35].
@@ 3.3 Training Strategy
Figure 4: Visual comparison of novel view synthesis on the D2RF defocus blur dataset [35]. Here, we also compare with methods fed with deblurred images produced by a state-of-the-art video deblurring method [70] to manifest the effectiveness of our method.
@@ 3.4 Loss Function
Figure 5: Visual comparison of novel view synthesis on the DyBluRF motion blur dataset [50].
@@ 5 Conclusion
Acknowledgement. This work was supported by the National Natural Science Foundation of China (62471499), the Guangdong Basic and Applied Basic Research Foundation (2023A1515030002).
@@ B.1 Novel View Synthesis Comparison
D2RF-v2 and DyBluRF-v2 Dataset. We evaluate our method on two datasets (DyBluRF-v2 and D2RF-v2) with both motion and defocus blur occurring simultaneously. Note that we obtain the DyBluRF-v2 dataset by applying depth-of-field (DoF) rendering technique in Bokehme [45] to the original DyBluRF dataset [50] to simulate defocus blur, and create the D2RF-v2 dataset with motion blur by processing the original D2RF dataset [35] using the motion blur generation method in Davanet [74]. Table 5 present the comparison results. As shown, our method clearly outperforms all the compared methods on the two datasets, verifying the advantage of our method in handling cases with motion and defocus blur occurring jointly.
Deblur-NeRF Dataset. We compare our approach against De3DGS [24], which is designed to reconstruct sharp static scenes from blurred static images, and evaluate them on the Deblur-NeRF [37] dataset. Table 6 and Figure 12 present the comparison results. Clearly, our method demonstrates significant advantages over other methods, producing sharper novel view images.
@@
Figure 10: Visual comparison of novel view synthesis on the DyBluRF motion blur dataset [50]. Here, we also compare with methods fed with deblurred images produced by a state-of-the-art video deblurring method [70] to manifest the effectiveness of our method.
@@
B.2 Deblurring Comparison
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
De3DGS [67] Ours GT
Figure 12: Visual comparison of novel view synthesis on the Deblur-NeRF dataset [37].
Table 5: Quantitative comparison of novel view synthesis on the D2RF-v2 defocus blur dataset and the DyBluRF-v2 motion blur dataset. The numerical results of defocus blur are obtained on the Shop and Car scenes of the D2RF-v2 dataset, and the numerical results of motion blur are obtained on the Man and Seesaw scenes of the DyBluRF-v2 dataset.
Method Defocus Blur Motion Blur
[…]
D3DGS [67] 23.66 0.739 0.257 21.75 0.655 0.289
SoM [55] 29.04 0.820 0.094 27.28 0.791 0.103
D2RF [35] 27.82 0.795 0.132 25.89 0.722 0.133
DyBluRF [50] 27.30 0.771 0.150 26.54 0.753 0.112
De4DGS [61] 29.74 0.856 0.078 27.97 0.824 0.087
Ours 30.26 0.885 0.062 28.55 0.859 0.064
Table 6: Quantitative comparison of novel view synthesis on the Deblur-NeRF dataset [37].
Method Defocus Blur Motion Blur
De3DGS [24] 23.71 0.747 0.110 26.61 0.822 0.108
Ours 24.22 0.768 0.095 27.14 0.835 0.096
@@ B.2 Deblurring Comparison
Appendix C Additional Ablation Results
C.1 Ablation on BP-Net
We conduct an ablation study to evaluate the contribution of BP-Net. Specifically, we compare three different blur modeling methods: (i) the motion blur and defocus blur modeling method used in De3DGS [24] (w/ blur modeling in De3DGS [24]), (ii) the motion blur modeling method in De4DGS [61] (w/ blur modeling in Deblur4DGS [61]), and (iii) the defocus blur modeling method in D2RF [35] (w/ blur modeling in D2RF [35]). We report the quantitative results in Table 8, where we can see that our method with the proposed BP-Net produces better results than these alternatives, demonstrating the effectiveness of the BP-Net.
Table 8: Effect of BP-Net.
@@ C.1 Ablation on BP-Net
w/ blur modeling in De3DGS [24] 28.31 0.812 0.098 26.05 0.823 0.118
w/ blur modeling in De4DGS [61] 28.63 0.829 0.094 26.74 0.859 0.060
w/ blur modeling in D2RF [35] 28.96 0.832 0.094 26.30 0.825 0.109
Ours with BP-Net 29.39 0.859 0.078 27.01 0.876 0.056
C.2 Ablation on Blur Kernel Size
Table 9 further perform quantitative evaluation on how different blur kernel sizes (denoted as $K$) affect the performance of our method. As shown, a larger blur kernel helps to obtain better results. However, this trend becomes less obvious when $K$ is larger than 9. To balance the performance and the computational cost, we thus choose $K=9$ as our default choice.
Table 9: Effect of varying blur kernel size $K$. The numerical results of defocus blur are obtained on the Gate and Dock scenes of the D2RF dataset, and the numerical results of motion blur are obtained on the Skating and Man scenes of the DyBluRF dataset.
@@ C.2 Ablation on Blur Kernel Size
Figure 14: Visualization of the blur kernel and blur intensity predicted by BP-Net. Note that the top row shows an image with defocus blur, while the bottom row shows an image with motion blur. In the blur ground truth (GT) image, blue markers indicate pixels with almost no blur, green markers denote pixels with mild blur, and red markers represent pixels with severe blur. A higher blur intensity value corresponds to a more heavily blurred area. Clearly, BP-Net can accurately predict blur regions in images with different types of blur and estimate the corresponding blur kernels at pixel locations with varying blur levels.
Figure 15: Visual comparison of estimated kernel on the D2RF-v3 defocus blur dataset and the DyBluRF-v3 motion blur dataset. Note that the top row shows an image with defocus blur, while the bottom row shows an image with motion blur. Clearly, BP-Net can accurately estimate blur kernels from various types of blurry images and thus recover sharp images.
Appendix D Robustness to Preprocessing and Segmentation Errors
Our method represents image blurring effects using two components: a blur kernel $k$ and a blur intensity $m$, as illustrated in Figure 14. The blur intensity $m$ effectively emphasizes the spatial regions affected by blur within each training image. Moreover, the type of blur can be intuitively inferred from the estimated kernels. Specifically, kernels corresponding to motion blur capture structured trajectories that reflect the camera’s movement, while those associated with defocus blur present Gaussian-shaped patterns that vary with the distance of the pixel from the focal plane.
To evaluate the accuracy of the blur kernel $k$ predicted by BP-Net, we compare the ground truth blur kernel with the blur kernel predicted by BP-Net on a blurry monocular video dataset with ground truth blur kernel. To this end, we construct two datasets with ground truth blur kernels, referred to as D2RF-v3 and DyBluRF-v3, by randomly sampling two global Gaussian and linear distribution blur kernels of size 9 x 9 and then respectively applying them to the ground truth sharp images in D2RF [35] and DyBluRF [50] to obtain the corresponding blurry images. With the two datasets, we quantitatively compare our estimated blur kernels and the ground truth blur kernels using PSNR and KL divergence as metrics. Table 10 and Figure 15 present the comparison results. Clearly, our estimated blur kernels are highly similar to the ground truth blur kernels in numerical metrics, manifesting the effectiveness of the BP-Net in predicting different types of blur.
Table 10: Comparison of ground truth kernels and estimated kernels.
D2RF-v3 DyBluRF-vs
[…]
Appendix E Robustness to Preprocessing and Segmentation Errors